OpenStudy (anonymous):

will give medal and fan best answerer Given the system of equations, what is the y-coordinate of the solution? (this is a matrix problem btw) 5x - 4y = 7 x = 5 - y 36/7 36/23 18/23

OpenStudy (anonymous):

In this situation, you can do two things: change both equations to equal to y and then set them equal to each other or use a scalar and subtract/add the equations. In this case, I like to use the scalar as fractions are really messy.

OpenStudy (anonymous):

Solve for x and then substitute the x-value into one of the equations(either is fine) to find the y-value.

OpenStudy (anonymous):

@Grazes you see this is what confuses me i understand that method but it gives you three correct answers for y witch one would it be?

OpenStudy (anonymous):

What. Hang on. I'll do it.

OpenStudy (anonymous):

I don't know what you're talking about but y should be 2 and there should only be one value of y that is true for an x in a system of equations.

OpenStudy (anonymous):

thanks for trying @Grazes

OpenStudy (anonymous):

@arilove1d can you please help

OpenStudy (anonymous):

y =2 STEP 1: Multiply second equation by -5. After multiplying we have the following system: 5x−4y−5x−5y=7=−25 STEP 2: add the two equations together to eliminate x from the system. −9y=−18 STEP 3: find y y=2 STEP 4: substitute the value for y into the original equation to solve for x. 5x−4(2)=7 x=3

OpenStudy (anonymous):

@arilove1d ya thats what @Grazes said but it gives me 3 answer for the y coordinate 36/7 36/3 18/23

OpenStudy (anonymous):

Hmmm i know its not the last one that would give you an decimal

OpenStudy (anonymous):

@arilove1d ya this confuses me majorly.... since when was math hard :/

OpenStudy (anonymous):

@csscgirl THIS CONFUSES ME!!! pleasehelp

OpenStudy (anonymous):

@csscgirl it is a matrix

OpenStudy (anonymous):

the y is 2, so idk sorry... but... just a sec.

OpenStudy (anonymous):

or 25/9... idk

OpenStudy (anonymous):

its okay thanks for your help @csscgirl i know ts super confusing...

OpenStudy (anonymous):

yw.... umm... @Mertsj

OpenStudy (anonymous):

@Mertsj do you understand

OpenStudy (mertsj):

\[5(5-y)-4y=7\] \[25-5y-4y=7\] \[25-9y=7\] \[-9y=-18\] \[y=2\]

OpenStudy (anonymous):

@mertj i literally copied this strait from my math website and pasted it and there are three option i know that y=2 and x=3 but i dont understand how t would equal any of the options... its called matrices the lesson im on if that helps

OpenStudy (mertsj):

Give me a link to the math website and let me look at it.

OpenStudy (mertsj):

That's only a login page.

OpenStudy (mertsj):

I can't see it unless you give me you user name and password.

OpenStudy (mertsj):

I don't think you got the list of answers that goes with that question.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

i hate this math class